OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = 6x^2 – 24x + 32?

OpenStudy (anonymous):

–2, –8) (–2, 8) (2, –8) (2, 8)

OpenStudy (anonymous):

I Need Someone To Show Me How To Do This Problem

OpenStudy (campbell_st):

ok... need to do some factoring y = 6(x^2 - 4x) + 32 can you complete the square for x^2 - 4x ..?

OpenStudy (anonymous):

@campbell_st I Think I Can Can You Just Show Me How To Do It So I Can Get A Sense Of How To Do The Other Problems

OpenStudy (campbell_st):

ok... so find half of the middle term (-4/2) then square it so you will have y = 6(x^2 - 4x + 4) + 32 ehich means your original problem is now y = 6x^2 _24x + 24 + 8 24 + 8 = 32 I just split the constant so factoring you have y = 6(x^2 - 4x + 4) + 8 or y = 6(x - 2)^2 + 8 the general form is y = (x - h)^2 + k the vertex is at (h, k) you need to identify the values of h and k... for the vertex in your question.